OpenStudy (anonymous):

prove n>equal to 5, 4n<2^n

OpenStudy (jhannybean):

since n \(\ge\) 5, we can plug 5 in for n. \(\large 4(5) < 2^5 \iff 20 < 32\)

OpenStudy (jhannybean):

If you think about it, \(\large 2^n\) is a function that is approaching \(+ \infty\) much,much faster than \(4n\), therefore all the numbers from 5 onwards are going to respectively represent the difference in these function's increase.

OpenStudy (anonymous):

its for a proofs class I need to show this through induction

OpenStudy (jhannybean):

Ohh I see....

OpenStudy (anonymous):

Okay, lets do some mathematical induction.

OpenStudy (anonymous):

Prove 4n < 2^n, for all n >= 5.

OpenStudy (anonymous):

Base case. n = 5. 4*5 = 2- <= 2^5 = 32 is true.

OpenStudy (anonymous):

Induction step. Assume 4n <= 2^n is true. Plug in (n+1) for n. 4(n+1) <= 2^(n+1). 4n + 4 <= 2*2^n 4n <= 2*2^n - 4 Now compare 2^n and 2*2^n - 4 (for n>=5). We can use induction a second time to prove that 2^n <= 2*2^n - 4 for all n>=5. Base case, n=5. 2^5 = 16 <= 2*2^5 - 4 = 28 is true. Induction step, Assume 2^n <= 2*2^n - 4 for all n>=5 is true. Plug (n+1) for n. 2^(n+1) <= 2*2^(n+1) - 4 2*2^n <= 4*2^n - 4 Divide both sides of the inequality by 2. 2^n <= 2*2^n -2 We know that 2*2^n - 4 <= 2*2^n - 2 is true, and since we assume 2^n <= 2*2^n - 4 we know the following is also true: 2^n <= 2*2^n - 4 <= 2*2^n - 2. This means that 2^n <= 2*2^n - 2 is true, which implies 2^(n+1) <= 2*2^(n+1) - 4 is true. By mathematical induction we conclude that 2^n <= 2*2^n - 4 is true for n>=5. Now, back to our original induction proof. Since we now know 2^n <= 2*2^n - 4, we can combine this with our assumption that 4n <= 2^n to get the following: 4n <= 2^n <= 2*2^n - 4. This means 4n < 2*2^n - 4, which implies 4(n+1) <= 2^(n+1) is true for n>=5. Therefore, by strong induction (using induction twice), we conclude 4n <= 2^n is true for all n >= 5.